I have $5.00 in nickels, dimes, and quarters. I have 3 times as many nickels as dimes. I have 5 times as much money in quarters as I do in nickels. How many of each coin do I have?
Answer by jorel1380(3719)

x+y+z=?
5x+10y+25z=500
x=3y
25/5=5 times the money so x=z
5(3y)+10y+25(3y)=500
15y+10y+75y=500
100y=500
y=5
x=15
z=15
You have 15 nickels, 5 dimes, and 15 quarters, for a total of $5.00 made up of 35 coins..